Grade 10IIT JEE Entrance Exam

in an a.p. the sum of `n` terms is 5n square-5n find tenth term

Pranjal Shinde
9 Years agoGrade 10

1 Answer

Jitendra Kumar
9 Years ago
Let $S1$ be the sum of the given AP $\rightarrow S1= 5n2-5n$
First terms is $a1=S1 = 0$ and $S2=a1+a2 =10=> a2 = 10$
So  common difference  $d = 10$
Thus general term is $an = a1+(n-1)d=10(n-1)$
Hence tenth term is $a10=10(10-1)=90$